# Clear Nights Ahead: Towards Multi-Weather Nighttime Image Restoration

### v1

@@ 6 Conclusion
This paper explores a practical yet under-explored task, i.e., multi-weather nighttime image restoration. To facilitate this task, we introduce an illumination-aware degradation generation approach and construct a new dataset featuring 10K high-quality nighttime images with various compositional adverse weather and lighting conditions. In addition, we propose ClearNight, a unified framework, tailored for the new task, capable of removing multiple degradations in one go. ClearNight leverages Retinex-based dual priors to explicitly guide the network to focus on uneven illumination regions and intrinsic texture contents respectively. Moreover, ClearNight incorporates weather-aware dynamic specific-commonality collaboration to better capture the characteristics of various weather conditions, enabling effective multi-weather degradation removal. Comprehensive experiments and analysis on both synthetic and real-world images demonstrate the superiority of our ClearNight.
[…]
@@
Appendix A Additional Implementation Details
The model was trained with a batch size of 1. To stabilize training, we adopt a cosine annealing warm restarts strategy with a restart period of 50 epochs. The proposed framework is trained and tested on an NVIDIA RTX A6000.
The architecture of ClearNight, illustrated in Figure 2, applies a multi-branch structure [60, 38, 68], integrating illumination and reflectance priors into the backbone. The backbone consists of a commonality branch and a weather-aware dynamic specificity branch. The commonality branch employs five Transformer blocks [43] with modified normalization layers, activation functions, and spatial feature aggregation to enhance global feature extraction for improved restoration performance. The weather-aware dynamic specificity branch includes five weather-aware dynamic selection blocks (WDS) tailored to adverse weather degradations. After each interaction between unique and common features, we use upsampling and downsampling operations to adjust feature map scales, doubling or halving the width and height as needed.
Appendix B Dataset License and Intended Use
@@
Appendix C Limitations and Future Works
@@
While our approach effectively handles nighttime illumination and various adverse weather degradations, its performance under extremely dynamic lighting conditions, such as rapidly changing artificial lighting, needs improvement. As shown in Figure 12, while most weather-induced degradations can be removed, rapidly changing flares still remain. Future work could focus on improving the robustness to such extreme scenarios. The original data for AllWeatherNight are sourced from BDD100K [56] and Exdark [36], which exclusively focus on driving and object detection. Although our dataset incorporates various nighttime adverse weather degradations, the natural scenes in it are relatively homogeneous, limiting its environmental diversity. We consider expanding the AllWeatherNight to include a wider range of real-world scenarios to further enhance the model’s generalization ability.
Appendix D Additional Analysis of AllWeatherNight
Analysis of AllWeatherNight’s Authenticity. To validate the authenticity of the AllWeatherNight dataset, we conduct a statistical analysis of feature distribution across synthetic adverse weather nighttime datasets and real-world samples. The synthetic datasets includes haze datasets (UNREAL-NH [33], GTA5 [52], NHC&NHR&NHM [64], YellowHaze&NightHaze [31]), rain datasets (GTAV-NightRain [59], GTAV-NightRain-H [60]) and snow dataset (RVSD [3]). As shown in Figure 13, the t-SNE visualization reveals that AllWeatherNight’s adverse weather degradation distributions are more closely aligned with real-world nighttime images. From left to right, the subfigures depict the feature distributions for haze, rain, and snow, respectively. Notably, AllWeatherNight’s haze distribution closely matches that of real-world nighttime haze, while its rain and snow distributions exhibit the highest similarity to real-world samples. These findings demonstrate that AllWeatherNight outperforms existing synthetic datasets in approximating real-world nighttime degradation features, making it highly suitable for training models that generalize to diverse real-world scenarios. Consequently, AllWeatherNight provides a robust foundation for enhancing model performance in multi-weather nighttime image restoration tasks.
(a) visualization of haze dataset features
(b) visualization of rain dataset features
(c) visualization of snow dataset features
@@ Appendix D Additional Analysis of AllWeatherNight
Figure 13: The t-SNE map of various nighttime adverse weather datasets. The feature distribution of adverse weather degradations in our AllWeatherNight is closer to real-world nighttime scenes than existing nighttime synthetic datasets.
@@
(a) without $\mathcal{L}_{\textit{bce}}$ and $\mathcal{L}_{\textit{lb}}$
(b) with only $\mathcal{L}_{\textit{bce}}$
(c) with only $\mathcal{L}_{\textit{lb}}$
(d) with both $\mathcal{L}_{\textit{bce}}$ and $\mathcal{L}_{\textit{lb}}$
Figure 15: Correlation between degradation types and candidate units correlation under different loss supervision.
Figure 16: Ablation study on $\lambda_{\textit{bce}}$.
Supplemental Visualizations of Illumination-Aware Generation. Existing synthetic adverse weather nighttime datasets typically model a single type of weather distortion and rarely account for background blurring induced by artificial lighting. In contrast, the proposed AllWeatherNight encompasses a diverse range of degradation, including haze, rain streaks, raindrops, snow, flare, and their combinations. Moreover, we design an illumination-aware degradation generation method to simulate realistic adverse weather and flare effects in nighttime scenes. To verify the effectiveness of the generation method, we synthesized datasets under four settings, with supplemental visualizations provided in Figure 14. Each type of degradation in AllWeatherNight varies in scales, directions, densities, and styles, ensuring comprehensive coverage. The generated flare preserves background clarity in the original image to some extent, enabling the network to effectively restore natural lighting and fine details.
Appendix E Additional Analysis of WDS
@@
Analysis of the Loss Functions. Our method employs the loss functions $\mathcal{L}_{\textit{lb}}$ and $\mathcal{L}_{\textit{bce}}$ to supervise the dynamic allocator and weather instructor, respectively, enabling effective auxiliary selection tasks. To evaluate their impact, we conduct experiments on AllWeatherNight, with results visualized in Figure 15. Evidently, $\mathcal{L}_{\textit{lb}}$ optimizes the allocation of candidate units, ensuring efficient utilization. Without $\mathcal{L}_{\textit{lb}}$, as shown in Figure 15, many units remain underutilized, degrading performance in complex multi-weather scenes. Similarly, applying $\mathcal{L}_{\textit{bce}}$ enhances the correlation between diverse weather conditions and candidate units selection, improving the model’s ability to address multiple adverse weather effects. Without $\mathcal{L}_{\textit{bce}}$, the selected units for different weather distortions tend to converge, impairing multi-weather nighttime image restoration, as shown in Figure 15. In contrast, the combined use of $\mathcal{L}_{\textit{bce}}$ and $\mathcal{L}_{\textit{lb}}$ enables precise identification of different adverse weather degradations and balanced utilization of candidate units. As depicted in Figure 15, the correlation visualization reveals that distinct units are selected for different adverse weather conditions, with similar sets of units consistently chosen for the same weather type, and all units are effectively utilized, maintaining balanced load distribution. Consequently, this dual-loss collaboration strategy significantly enhances the model’s performance in restoring complex nighttime scenarios with diverse adverse weather conditions and flare effects, achieving high-quality results.
@@
The weighting of the loss function critically influences the performance of the network during training, particularly for multi-weather nighttime scenes with flare effects. To determine an optimal weighting stratregy for $\mathcal{L}_{\textit{bce}}$, we conduct an ablation study on the Rain Scene in AllWeatherNight dataset, with results shown in Figure 16. The absence of $\mathcal{L}_{\textit{bce}}$ significantly degrades model’s performance, as it fails to distinguish diverse adverse weather degradation features, impairing the guidance for unit selection. When the weight was set to 0.01, the overly strong constraint of the $\mathcal{L}_{\textit{bce}}$ limits its restoration performance for different weather with similar patterns. Conversely, with an excessively low weight of 0.0001, various adverse weather features cannot be effectively distinguished, resulting in poor performance across multiple weather conditions. Weights of 0.001 achieve the best performance. Therefore, we set the $\lambda_{\textit{bce}}$ to 0.001 for our model.
Appendix F Analysis of Computational Complexity
@@
Figure 17 compares the parameters and inference time of our method with recent competitive approaches. While diffusion-based method such as WeatherDiff [69] and DiffUIR [65] achieve impressive performance, they suffer from higher inference time, which limits their practicality in real-world scenes. On the other hand, lightweight models such as RAMiT [8] and DEA-Net [7] demonstrate fast inference, but often compromise restoration quality. Our ClearNight achieves a good trade-off between performance and efficiency. With only 2.84M parameters and an inference time of 0.32s, it is significantly faster than most diffusion-based models while offering much stronger restoration performance than lightweight models. In summary, ClearNight maintains a relatively low computational cost, making it a practical and effective solution for multi-weather nighttime image restoration.
Appendix G Additional Visualization Results
@@
Multi-Degradation. We have provided more visual results of our model on Rain Scene and Snow Scene in AllWeatherNight, as shown in Figure 18. The Rain and Snow Scene testing subsets contain degraded images with complex multi-weather degradations, effectively simulating real-world nighttime scenes. The visual results demonstrate that ClearNight excels in removing the influence of adverse weather while restoring rich background details. Especially when the background is similar to the degradation caused by the entanglement of adverse weather and flares, ClearNight accurately distinguishes these factors, producing high-quality images.
@@
Single-Degradation. Although multiple adverse weather conditions frequently occur in an intertwined manner, the restoration of simple weather nighttime scenes should not be ignored. We also evaluate the generalization capability of ClearNight on single-degradation conditions. As shown in Figure 19, we can see that our method is equally effective in single-degradation scenarios.
@@
Real-World. Nighttime real-world degraded images often contain complex weather interferences and diverse flares. As shown in Figure 20, we provide more comparison result of real-world scenes, e.g., haze ($1^{\textit{st}}$ and $2^{\textit{nd}}$ rows), rain ($3^{\textit{rd}}$ and $4^{\textit{th}}$ rows), raindrop ($5^{\textit{th}}$ and $6^{\textit{th}}$ rows), snow ($7^{\textit{th}}$ and $8^{\textit{th}}$ rows). The rain and snow scenes often exhibit subtle haze effects, as evidenced by real-world nighttime images, consistent with multi-degradation scenarios in our dataset. The restored results show that ClearNight effectively removes the interference of multiple adverse weather conditions while recovering natural nighttime lighting. Compared to existing adverse weather image restoration approaches, our method is superior in reconstructing scene content and lighting conditions.

### v2

@@ Conclusion
This paper explores a practical yet under-explored task, i.e., multi-weather nighttime image restoration. To facilitate this task, we introduce an illumination-aware degradation generation approach and construct a new dataset featuring 10K high-quality nighttime images with various compositional adverse weather and lighting conditions. In addition, we propose ClearNight, a unified framework, tailored for the new task, capable of removing multiple degradations in one go. ClearNight leverages Retinex-based dual priors to explicitly guide the network to focus on illumination regions and intrinsic textures respectively. Moreover, ClearNight incorporates weather-aware dynamic specific-commonality collaboration to better capture the characteristics of diverse weather, enabling effective multi-weather degradation removal. Comprehensive experiments on both synthetic and real-world images demonstrate the superiority of our ClearNight.
Clear Nights Ahead: Towards Multi-Weather Nighttime Image Restoration
@@
Additional Implementation Details
As detailed in the main text, the model was trained on 256 $\times$ 256 randomly cropped patches. For inference, full-resolution images were used. The model was trained with a batch size of 1. To stabilize training, we adopt a cosine annealing warm restarts strategy with a restart period of 50 epochs. We use the Adam optimizer (Adam) with $\beta_{1}$=0.9 and $\beta_{2}$=0.99. The proposed framework is implemented in Pytorch and was trained and tested on an NVIDIA RTX A6000.
The architecture of ClearNight, applies a multi-branch structure (NICCV2023; UCVPR2023; LWGWSCVPR2023), integrating illumination and reflectance priors into the backbone. The backbone consists of a commonality branch and a weather-aware dynamic specificity branch. The commonality branch employs five-stage Transformer blocks (swin) with modified normalization layers, activation functions, and spatial feature aggregation to enhance global feature extraction for improved restoration performance. The weather-aware dynamic specificity branch includes five weather-aware dynamic selection blocks (WDS) tailored to adverse weather degradations. After each interaction between unique and common features, we use upsampling and downsampling operations to adjust feature map scales, doubling or halving the width and height as needed.
General Discussions
Dataset License and Intended Use
@@
Limitations and Future Works
@@
Figure 13: The t-SNE map of various nighttime adverse weather datasets, including haze, rain and snow degradations. The feature distribution of adverse weather degradations in our AllWeatherNight is closer to real-world nighttime scenes than existing nighttime synthetic datasets.
Figure 14: Models trained with AllWeatherNight achieve superior performance on real-world images compared to those trained with the combination of existing nighttime adverse weather datasets in Tab. 1.
While our approach effectively handles nighttime illumination and various adverse weather degradations, its performance under extremely dynamic lighting conditions, such as rapidly changing artificial lighting, needs improvement. As shown in Fig. 12, while most weather-induced degradations can be removed, rapidly changing flares still remain. Future work could focus on improving the robustness to such extreme scenarios. The original data for AllWeatherNight are sourced from BDD100K (BDD100K) and Exdark (Exdark), which exclusively focus on driving and object detection. Although our dataset incorporates various nighttime adverse weather degradations, the natural scenes in it are relatively homogeneous, limiting its environmental diversity. We consider expanding the AllWeatherNight to include a wider range of real-world scenarios to further enhance the model’s generalization ability.
Further Analysis of AllWeatherNight Dataset
Authenticity Assessment
To validate the authenticity of the AllWeatherNight dataset, we conduct a statistical analysis of feature distribution, comparing it against existing synthetic adverse weather nighttime datasets and real-world samples. The nighttime synthetic datasets include various haze datasets (UNREAL-NH (NACM20232), GTA5 (NECCV2020), NHC/NHR/NHM (NACM2020), NightHaze/YellowHaze (NPCM2018)), rain datasets (GTAV-NightRain (GTAV), GTAV-NightRain-H (NICCV2023)) and snow dataset (RVSD (SICCV2023)). As shown in Fig. 13, the t-SNE visualization reveals that AllWeatherNight’s adverse weather degradation distributions are more closely aligned with real-world nighttime images than any other synthetic nighttime dataset. From left to right, the subfigures depict the feature distributions for haze, rain, and snow, respectively. Notably, AllWeatherNight’s haze distribution closely matches that of real-world nighttime haze, while its rain and snow distributions exhibit the highest similarity to real-world adverse weather samples. These findings demonstrate that AllWeatherNight outperforms existing synthetic datasets in approximating real-world nighttime degradation features, making it highly suitable for training models that generalize to diverse real-world scenarios. Consequently, AllWeatherNight provides a robust foundation for enhancing model performance in multi-weather nighttime image restoration tasks.
Effectiveness on Real-World Data
To validate the effectiveness of our AllWeatherNight, we train baseline and ClearNight models on our dataset and on a composite dataset, which includes all existing nighttime single weather datasets listed in Tab. 1 of the main text. As shown in Fig. 14, both the baseline and our ClearNight models trained on AllWeatherNight consistently produce high-quality results on real-world nighttime scenarios. In contrast, models trained on the composite dataset tend to generate artifacts and often leave adverse weather degradations in the restored results. This underperformance can be attributed to the composite dataset’s fundamental failure to account for the co-occurrence of multiple weather conditions and the complex entanglement of these degradations with uneven lighting, consequently lacking real-world fidelity and realism. Consistent with quantitative results in Fig. 7 of the main text, these visual results demonstrate the superior quality and effectiveness of our AllWeatherNight for multi-weather nighttime image restoration.
@@
Figure 16: ClearNight outperforms a cascade of two latest task-specific nighttime adverse weather image restoration methods. Quantitative results are averaged over all the real-world rain streak images.
[…]
Figure 18: Ablation study on $\lambda_{\mathrm{b}}$.
Additional Visualizations of Illumination-Aware Degradation Generation
Existing synthetic adverse weather nighttime datasets typically model a single type of weather distortion and rarely account for background blurring induced by artificial lighting. In contrast, the proposed AllWeatherNight encompasses a diverse range of degradation, including haze, rain streaks, raindrops, snow, flare, and their combinations. Moreover, we design an illumination-aware degradation generation method to simulate realistic adverse weather and flare effects in nighttime scenes. We synthesized datasets under four settings, with supplemental visualizations provided in Fig. 15. Each type of degradation in AllWeatherNight varies in scales, directions, densities, and styles, ensuring comprehensive coverage. To simulate these degradations, we apply the atmospheric scattering model and interpolated Gaussian noise for haze and rain streaks (HCVPR2019), utilize Bezier curves for raindrop synthesis (Raindrop), and add snow degradation using snowflake masks from (snow100k; csd). The generated flare preserves background clarity in the original image to some extent, enabling the network to effectively restore natural lighting and fine details.
Further Analysis of ClearNight
ClearNight vs. Cascaded Solutions
To further validate the necessity of our unified framework, ClearNight, for restoring complex nighttime images under multiple adverse weather conditions, we present comprehensive analysis comparing our unified model against two pre-trained task-specific models (NICCV2023; NAAAI2025) and their cascaded pipeline. For the cascaded pipeline, we evaluate two configurations: one that applies the dehazing model followed by the deraining model (dehazing + deraining) and the other in the reverse order (deraining + dehazing). As illustrated in Fig. 16, task-specific models fail to remove multiple weather degradations, and their cascaded models struggle to produce satisfactory results. In contrast, our ClearNight generates more visually pleasing results and achieves the best average NIQE scores with superior computational efficiency.
Additional Ablation Studies
Influence of Retinex-Based Dual Prior Guidance. To evaluate the impact of multi-scale dual prior, we conduct ablation studies on illumination and reflectance priors at various scales, using # 1 and # 3 (in Tab. 4 of the main text) as baselines, respectively. As shown in Tabs. 5-6, incorporating multi-scale illumination and reflectance priors into the network enables it to focus on uneven lighting and intrinsic texture features of different scales, significantly enhancing nighttime image restoration performance. As the inclusion of illumination components across all five stages brings only marginal gains (PSNR + 0.1628, SSIM + 0.0022), we adopt them only in the first three stages to achieve a better trade-off between performance and model complexity.
Effectiveness of DSM. We investigate the number of candidate units ($B$) and top-$K$ selection parameter ($K$). As shown in Fig. 19, increasing the number of candidate units enhances selective capacity, with an optimal $K$ being critical for superior performance. Notably, performance improves with larger $K$ up to a threshold, beyond which irrelevant information causes model degradation.
Table 5: Ablation on Retinex-based Illumination Prior.
[…]
1 28.7976 / 0.8825
2 31.6843 / 0.9109
3 31.6714 / 0.9114
4 31.6429 / 0.9110
5 31.7217 / 0.9115
6 32.1304 / 0.9176
Table 6: Ablation on Retinex-based Reflection Prior.
[…]
1 31.7075 / 0.9113
2 31.9680 / 0.9198
3 31.9374 / 0.9201
4 31.8438 / 0.9172
5 32.0335 / 0.9203
6 32.0594 / 0.9205
@@
Analysis of the Loss Functions. Our method employs the loss functions $\mathcal{L}_{\mathrm{lb}}$ and $\mathcal{L}_{\mathrm{b}}$ to supervise the dynamic allocator and weather instructor, respectively, enabling effective auxiliary selection tasks. To evaluate their impact, we conduct experiments on AllWeatherNight, with results visualized in Fig. 17. Evidently, $\mathcal{L}_{\mathrm{lb}}$ optimizes the allocation of candidate units, ensuring efficient utilization. Without $\mathcal{L}_{\mathrm{lb}}$, many units remain underutilized, degrading performance in complex multi-weather scenes. Similarly, applying $\mathcal{L}_{\mathrm{b}}$ enhances the correlation between diverse weather conditions and candidate units selection, improving the model’s ability to address multiple adverse weather effects. Without $\mathcal{L}_{\mathrm{b}}$, the selected units for different weather distortions tend to converge, impairing multi-weather nighttime image restoration. In contrast, the combined use of $\mathcal{L}_{\mathrm{b}}$ and $\mathcal{L}_{\mathrm{lb}}$ enables precise identification of different adverse weather degradations and balanced utilization of candidate units.
As depicted in the bottom-right of Fig. 17, the correlation visualization reveals that distinct units are selected for different adverse weather conditions, with similar sets of units consistently chosen for the same weather type, and all units are effectively utilized, maintaining balanced load distribution. Consequently, this dual-loss collaboration strategy significantly enhances the model’s performance in restoring complex nighttime scenarios with diverse adverse weather conditions and flare effects, achieving high-quality results.
@@
The weighting of the loss function critically influences the performance of the network during training, particularly for multi-weather nighttime scenes with flare effects. To determine an optimal weighting strategy for $\mathcal{L}_{\mathrm{b}}$, we conduct an ablation study on the Rain Scene in AllWeatherNight dataset, with results shown in Fig. 18. The absence of $\mathcal{L}_{\mathrm{b}}$ significantly degrades model’s performance, as it fails to distinguish diverse adverse weather degradation features, impairing the guidance for unit selection. When the weight was set to 0.01, the excessively strong constraint of $\mathcal{L}_{\mathrm{b}}$ limits its restoration performance on different adverse weather degradations with similar patterns. Conversely, with an excessively low weight of 0.0001, various adverse weather features cannot be effectively distinguished, resulting in poor performance across multiple weather conditions. Weights of 0.001 achieve the best performance. Therefore, we set the $\lambda_{\mathrm{b}}$ to 0.001 for our model.
Inference Efficiency
@@
Fig. 20 compares the average parameters and inference time of our method with recent competitive approaches across all synthetic scenes. While diffusion-based method such as WeatherDiff (WeatherDiff2023) and DiffUIR (SCVPR2024) achieve impressive performance, they suffer from higher inference time, which limits their practicality in real-world scenes. On the other hand, lightweight models such as RAMiT (RAMITCVPR2024) and DEA-Net (TIP2024DEA) demonstrate fast inference, but often compromise restoration quality. Our ClearNight achieves a good trade-off between performance and efficiency. With only 2.84M parameters and an inference time of 0.32s, it is significantly faster than most diffusion-based models while offering much stronger restoration performance than lightweight models. In summary, ClearNight maintains a relatively low computational cost, making it a practical and effective solution for multi-weather nighttime image restoration.
Additional Qualitative Comparisons
@@
Multi-Degradation. We have provided more visual results of our model on Rain Scene and Snow Scene in AllWeatherNight, as shown in Fig. 21. The Rain and Snow Scene testing subsets contain degraded images with complex multi-weather degradations, effectively simulating real-world nighttime scenes. The visual results demonstrate that ClearNight excels in removing the influence of adverse weather while restoring rich background details. Especially when the background is similar to the adverse weather degradation, ClearNight accurately distinguishes these factors, producing high-quality images.
@@
Single-Degradation. Although multiple adverse weather conditions frequently occur in an intertwined manner, the challenge of restoring scenes under simple degradations remains a crucial and non-trivial aspect of our work. Thus, we evaluate the generalization capability of ClearNight on single-degradation conditions. As shown in Fig. 22, we can see that our method is equally effective in single-degradation scenarios.
@@
Real-World. Nighttime real-world degraded images often contain complex weather interferences and diverse flares. As shown in Fig. 23, we provide more comparison result of real-world scenes, e.g., haze ($1^{\mathrm{st}}$ and $2^{\mathrm{nd}}$ rows), rain ($3^{\mathrm{rd}}$ and $4^{\mathrm{th}}$ rows), raindrop ($5^{\mathrm{th}}$ and $6^{\mathrm{th}}$ rows), snow ($7^{\mathrm{th}}$ and $8^{\mathrm{th}}$ rows). The rain and snow scenes often exhibit subtle haze effects, as evidenced by real-world nighttime images, consistent with multi-degradation scenarios in our dataset. The restored results show that ClearNight effectively removes the interference of multiple weather conditions while recovering natural nighttime lighting. Compared to existing adverse weather image restoration approaches, our method is superior in reconstructing scene content and natural lighting conditions.